Circular No. 5903

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505

MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET KUSHIDA-MURAMATSU (1993t)

S. Nakano, Sumoto, Japan, reports the photographic discovery of a comet by Yoshio Kushida and Osamu Muramatsu. The following observations have been reported:

1993 UT  α_{2000}  δ_{2000}  m_{1} Observer
Dec. 8.65382 5^{h}35^{m}49.^{s}02 +22^{o}33′05′′.1 16.5 Kushida
8.68646 5 35 47.44 +22 33 05.7  _{”}
9.54254 5 35 06.94 +22 33 25.7 16.5  _{”}
9.57813 5 35 05.26 +22 33 27.6  _{”}
11.81180 5 33 17.35 +22 34 19.6 16  _{”}
12.44468 5 32 46.96 +22 34 36.4 16.0 Kojima
12.45816 5 32 46.26 +22 34 37.6  _{”}
12.46795 5 32 45.81 +22 34 36.7 16  Urata
12.47517 5 32 45.46 +22 34 37.2  _{”}
12.47845 5 32 45.33 +22 34 37.1  _{”}
12.48333 5 32 45.14 +22 34 37.1 15.7 Nakamura
12.48681 5 32 44.97 +22 34 37.2  _{”}
12.48891 5 32 44.89 +22 34 37.0  Kobayashi
12.51372 5 32 43.63 +22 34 38.1  Nakamura
12.53315 5 32 42.60 +22 34 38.1  Kobayashi
12.53843 5 32 42.26 +22 34 37.3  Kojima
12.54076 5 32 42.23 +22 34 38.2  Kobayashi

Y. Kushida and O. Muramatsu (Yatsugatake South Base Observatory). 0.25-m f/3.4 reflector. Comet diffuse with central condensation; coma diameter ∼ 1′-2′.

T. Kojima (YGCO Chiyoda Observatory). 0.25-m Wright-Schmidt Camera

+ CCD. Coma extends to the west. T. Urata (Nihondaira Observatory Oohira Station). 0.31-m reflector +

CCD. Comet diffuse with moderate central condensation; coma diameter 30′′, extending toward p.a. 270^{o}.

A. Nakamura (Kuma Kogen). 0.60-m f/6.0 Ritchey-Chrétien + CCD (no

filter). Short tail toward p.a. 260^{o}. T. Kobayashi (Oizumi). 0.25-m f/4.4 reflector + CCD.

1993 December 12 Daniel W. E. Green